OpenStudy (anonymous):

What is the equation, in standard form, of a parabola that models the values in the table?

OpenStudy (anonymous):

x -2 , 0 , 4 f(x) -7 , 3 , -73

OpenStudy (anonymous):

A. y = -3x^2 - 4x + 3 B. y = -4x^2 - 3x + 3 C. y = -3x^2 - 3x + 4 D. y = 4x^2 + 3x - 3

OpenStudy (anonymous):

@ranga @ranga @ranga @ranga @ranga

OpenStudy (ranga):

When x = 0, f(x) = 3. Try that in each of the answer choices and you can quickly rule out two of the answers. Then, of the remaining two choices, try x = -2 and see which one gives you -7.

OpenStudy (anonymous):

How does that rule out 2 choices. which 2 choices are you talking about ?

OpenStudy (anonymous):

B and D are eliminated ?

OpenStudy (ranga):

When x = 0, A: gives 3 B: gives 3 C gives 4 D gives -3 Only the first two gives 3. The last two choices don't give 3 and so they are out.

OpenStudy (anonymous):

Okay C amd D are eliminated

OpenStudy (ranga):

Yes. Put x = -2 in the first two choices and see which one gives -7.

OpenStudy (anonymous):

neither one i got 23 and 25

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

A. y = -3x^2 - 4x + 3 x = -2 y = -3(-2)^2 - 4(-2) + 3 = -3(4) + 8 + 3 = -12 + 11 = -1 B. y = -4x^2 - 3x + 3 x = -2 y = -4(-2)^2 - 3(-2) + 3 = -4(4) + 6 + 3 = -16 + 9 = -7

OpenStudy (anonymous):

okay B is the answer .. lol ^_^

OpenStudy (anonymous):

Okay and one more quick question here ...

OpenStudy (anonymous):

\[\sqrt{-144}\] Is that 12i or 144i ?

OpenStudy (ranga):

12i

OpenStudy (anonymous):

Okay thanks again

OpenStudy (ranga):

You are welcome.